Hi I was wondering if someone could help me solve these 3 simultaneous equations with 3 unknown variables.
0.2857A - 0.0306B - 0.2857C = 0
-0.4286A - 0.2857B + 0.4286C = 0
-0.8571A - 0.8571B - 0.8571C + 1962 = 0
Thanks in advance.

Answer by richwmiller(17219)

0.2857A - 0.0306B - 0.2857C = 0
0.2857A=+ 0.0306B +0.2857C
A=0.0306B/0.2857 +C
A = 0.107105 B+C
-0.4286A - 0.2857B + 0.4286C = 0
.4286/0.2857(C-A) = B
1.50018 (C-A) = B
1.50018 (C-(0.107105 B+C)) = B
-0.160677B = B
B must be equal to 0
A=0.0306B/0.2857 +C
A=C
Remove B and substitute A for C
-0.8571A - 0.8571B - 0.8571C + 1962 = 0
-0.8571A - 0.8571A + 1962 = 0
A=1144.56
B=0
C=1144.56